# Geodesic Slice Sampler for Multimodal Distributions with Strong Curvature

### v1

@@ B.2 Numerical Integrators
model metr Wass min ESS avg ESS avg step-out avg shrinkage t(s)
aus euclidean [0.74, 0.12] [18, 5] [228, 12] [0.14, 0.0] [3.33, 0.01] 8.2
fisher [0.58, 0.01] [177, 9] [270, 13] [0.95, 0.0] [1.0, 0.01] 3298.6
ger euclidean [0.49, 0.01] [35, 10] [119, 9] [0.08, 0.0] [4.19, 0.02] 15.8
monge [0.75, 0.02] [80, 5] [4673, 113] [3.97, 0.03] [0.28, 0.01] 4665.4
fisher [0.49, 0.0] [85, 24] [169, 5] [0.95, 0.0] [0.99, 0.01] 19684.4
hrt euclidean [0.63, 0.01] [137, 26] [254, 16] [0.19, 0.01] [2.72, 0.02] 7.3
monge [0.74, 0.03] [394, 55] [2979, 266] [2.99, 0.04] [0.38, 0.02] 637.4
fisher [0.64, 0.01] [232, 14] [311, 8] [0.92, 0.01] [1.01, 0.01] 1694.8
pim euclidean [0.21, 0.0] [266, 41] [445, 45] [0.11, 0.0] [3.59, 0.03] 6.8
monge [0.29, 0.01] [515, 72] [4656, 269] [2.38, 0.03] [0.53, 0.02] 1711.7
fisher [0.21, 0.0] [427, 47] [547, 30] [0.93, 0.0] [0.98, 0.01] 425.4
rip euclidean [0.09, 0.01] [829, 112] [1499, 56] [0.24, 0.01] [2.46, 0.02] 4.2
monge [0.15, 0.02] [1042, 211] [3123, 685] [1.38, 0.02] [0.97, 0.01] 593.9
fisher [0.09, 0.0] [1623, 113] [1753, 92] [0.91, 0.0] [0.96, 0.01] 204.7
Table 1: Bayesian Logistic Regression, the notation is $[\mathrm{mean},\mathrm{std}]$.
@@ Appendix C Mathematical Derivations
C.1 The Generative Metric
The Generative metric components are:
$\operatorname{\boldsymbol{G}}(x)=f(x)\operatorname{\boldsymbol{I}}=\exp(\log f(x))\operatorname{\boldsymbol{I}},$
where the scalar factor is $f(x)=\left(\frac{p(x)+\lambda}{p_{0}+\lambda}\right)^{2}$.
@@ C.1 The Generative Metric
$G^{\tfrac{1}{2}}(x)$ $=\exp{\tfrac{1}{2}\log f(x)}\operatorname{\boldsymbol{I}},$
$G^{-\tfrac{1}{2}}(x)$ $=\exp{-\tfrac{1}{2}\log f(x)}\operatorname{\boldsymbol{I}},$
$\log|\det\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})|$ $=D\log f(x),$
where $f(x)$ is the scalar factor.
@@
Given the Riemannian metric
$G(x)=\left(\frac{p(x)+\lambda}{p_{0}+\lambda}\right)^{2}I,$
The metric tensor entries are
@@
The partial derivative of $f(x)$ with respect to $x_{i}$ is
$\partial_{i}f(x)=\frac{d}{dx_{i}}\left(\frac{p(x)+\lambda}{p_{0}+\lambda}\right)^{2}=2\left(\frac{p(x)+\lambda}{p_{0}+\lambda}\right)\frac{1}{p_{0}+\lambda}\frac{\partial p}{\partial x_{i}}.$
The Christoffel symbols for this metric are given by
@@
In matrix notation, let $e_{k}$ be the standard basis vectors,
$\Gamma^{k}=\frac{1}{2}\left(\nabla\log f(x)\,e_{k}^{\top}+e_{k}\,\nabla\log f(x)^{\top}-\partial_{k}\log f(x)\operatorname{\boldsymbol{I}}\right).$
The geodesic equations require computing
@@
The geodesic equations read
@@
C.2 The Monge metric
The Monge metric and its inverse,
$\operatorname{\boldsymbol{G}}(x)=\operatorname{\boldsymbol{I}}+\alpha^{2}\nabla\ell\nabla\ell^{\top},\quad\operatorname{\boldsymbol{G}}^{-1}(x)=\operatorname{\boldsymbol{I}}-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell}^{2}}\nabla\ell\nabla\ell^{\top}.$
@@ C.2 The Monge metric
Define the quantity $L_{\alpha}:=1+\alpha^{2}\norm{\nabla\ell}^{2}$, then the values are:
$G^{1/2}(x)$ $=\operatorname{\boldsymbol{I}}+\frac{\alpha^{2}}{1+\sqrt{L_{\alpha}}}\nabla\ell(\operatorname{\boldsymbol{x}})\nabla\ell(\operatorname{\boldsymbol{x}})^{\top},$
$G^{-1/2}(x)$ $=\operatorname{\boldsymbol{I}}-\frac{\alpha^{2}}{L_{\alpha}+\sqrt{L_{\alpha}}}\nabla\ell(x)\nabla\ell(x)^{\top},$
$\log|\det\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})|$ $=\log(1+\alpha^{2}\norm{\nabla\ell}^{2}).$
We now derive the quantities. Hartmann et al. [2022] gave:
$\operatorname{\boldsymbol{G}}^{-\tfrac{1}{2}}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{I}}+\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1}{\sqrt{L_{\alpha}}}-1\right)\nabla\ell\nabla\ell^{\top}$
Let us find a numerical sable form for $\operatorname{\boldsymbol{G}}^{-\tfrac{1}{2}}(\operatorname{\boldsymbol{x}})$. Multiplying by a convenient quantity, gives:
@@
$\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1}{\sqrt{L_{\alpha}}-1}\right)=\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1-\sqrt{L_{\alpha}}}{\sqrt{L_{\alpha}}}\right)\left(\frac{1+\sqrt{L_{\alpha}}}{1+\sqrt{L_{\alpha}}}\right)=\frac{\alpha^{2}}{L_{\alpha}+\sqrt{L_{\alpha}}}.$
The computation of $\operatorname{\boldsymbol{G}}^{\tfrac{1}{2}}(\operatorname{\boldsymbol{x}})$. For convenience take $\operatorname{\boldsymbol{y}}=\nabla\ell(x)$. The the metric is $\operatorname{\boldsymbol{G}}(y)=I+yy^{\top}$. Let us assume the square root is of the form $\operatorname{\boldsymbol{G}}^{\tfrac{1}{2}}(\operatorname{\boldsymbol{y}})=\operatorname{\boldsymbol{I}}+\lambda\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}$. Let us formulate a quadratic equation for $\lambda$:
$\operatorname{\boldsymbol{G}}^{\tfrac{1}{2}}(\operatorname{\boldsymbol{y}})\operatorname{\boldsymbol{G}}^{\tfrac{1}{2}}(\operatorname{\boldsymbol{y}})$ $=\operatorname{\boldsymbol{I}}+\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}$
$I+2\lambda\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}+\lambda^{2}\norm{\operatorname{\boldsymbol{y}}}^{2}\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}$ $=\operatorname{\boldsymbol{I}}+\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}$
00 0 $=\left(1-2\lambda-\lambda^{2}\norm{\operatorname{\boldsymbol{y}}}^{2}\right)\operatorname{\boldsymbol{y}}\operatorname{\boldsymbol{y}}^{\top}.$
The solutions of the quadratic equation are:
$\lambda=\frac{-1\pm\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}}{\|\operatorname{\boldsymbol{y}}\|^{2}}.$
Let us simplify $\frac{-1+\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}}{\|\operatorname{\boldsymbol{y}}\|^{2}}$, multiply by conjugate
$\frac{\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}-1}{\|\operatorname{\boldsymbol{y}}\|^{2}}\left(\frac{\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}+1}{\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}+1}\right)=\frac{\norm{\operatorname{\boldsymbol{y}}}^{2}}{\norm{\operatorname{\boldsymbol{y}}}^{2}\sqrt{1+\|\operatorname{\boldsymbol{y}}\|^{2}}+1}.$
Substitute $\operatorname{\boldsymbol{y}}=\alpha\nabla\ell(\operatorname{\boldsymbol{x}})$ and we obtain the result:
@@
$G^{1/2}(\operatorname{\boldsymbol{x}})=\operatorname{\boldsymbol{I}}+\frac{\alpha^{2}}{\sqrt{1+\alpha^{2}\|\nabla\ell(\operatorname{\boldsymbol{x}})\|^{2}}+1}\nabla\ell\nabla\ell^{\top}.$
Christoffel symbols
For the Monge metric the Christoffel symbols taken from Hartmann et al. [2022] are:
$\Gamma^{k}(\operatorname{\boldsymbol{x}})$ $=\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell}^{2}}\nabla^{2}\ell\partial_{k}\ell.$
The geodesic equations read
@@
The Christoffel symbols associated to the inverse Monge metric are, the derivation in Section C.4:
$\Gamma^{k}=\frac{\alpha^{2}}{2}\Big{[}L_{\alpha}\left(\nabla f\nabla\ell^{\top}+\nabla\ell\nabla f^{\top}+2f\nabla^{2}\ell\right)\partial_{k}\ell+\nabla\ell\nabla\ell^{\top}\partial_{k}f+\alpha^{2}\left\langle\nabla\ell,\nabla f\right\rangle\nabla\ell\nabla\ell^{\top}\partial_{k}\ell\Big{]}.$
The geodesic equations read:
@@ C.3 Monge metric: Christoffel symbols derivation
For completeness let us do a different derivation from Hartmann et al. [2022] of the Christoffel symbols. Take the auxiliary function $f(x)=-\frac{1}{L_{\alpha}}$, where $L_{\alpha}=1+\alpha^{2}\norm{\nabla\ell}^{2}$. The metric and inverse components are:
@@
Let us compute $\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}}$, which appears in the geodesic equations:
@@ C.4 Inverse Monge metric: Christoffel symbols derivation
Again the auxilary function is $f(x)=-\frac{1}{L_{\alpha}}$, where $L_{\alpha}=1+\alpha^{2}\norm{\nabla\ell}^{2}$, the metric and inverse components are:
@@
The derivatives of the metric are (we mark with the same color repeating terms):
@@
Let us compute the Christoffel symbols of the first kind (blue and red terms will cancel out, pink terms add to each other):
@@
The Christoffel symbols of the second kind read:
@@
Written in matrix form:
@@
Let us compute $\norm{v}^{2}_{\Gamma^{k}}$ which appears in the geodesic equations:
@@
Then the geodesic equations read:
@@
C.5 Complex geometry unimodal distributions
@@
$p(x)=\mathcal{N}(x_{D}|0,\sigma^{2})\mathcal{N}(x_{1:{D-1}}|\mu,e^{x_{D}}I_{D-1})$, use the notation $x=x(z)$ and $z=z(x)$, where $z\sim\operatorname{\mathcal{N}}(0,I)$. Choice of parameters $\sigma=3$, $\mu=0$.
@@
For simplicity here we show the two dimensional case, the full distribution can be consulted in Pagani et al. [2022]. The two dimensional density is: $p(x)=\operatorname{\mathcal{N}}(x_{1}|a,\frac{1}{2})\operatorname{\mathcal{N}}(x_{2}|x_{1}^{2},\frac{1}{2b})$. Choice of parameters $a=1$, $b=100$ and block size of $3$ and $\lfloor\tfrac{D-1}{3}\rfloor$ total blocks.
@@
The density is $p(x)=\operatorname{\mathcal{N}}(x(z)|\mu,\Sigma)|\det\partialderivative{x}{z}|$, where $z\sim\operatorname{\mathcal{N}}(\mu,\Sigma)$. Choice of parameters $a=1.5$, $\mu=0$, $\Sigma=\operatorname{\mathrm{diag}}(5,\tfrac{1}{2},..,\tfrac{1}{2})$
@@
For these toy problems we can express the Fisher Information follows from the transformation rule for Riemannian metrics:
$\operatorname{\boldsymbol{G}}(\operatorname{\boldsymbol{x}})=\partialderivative{\operatorname{\boldsymbol{z}}}{\operatorname{\boldsymbol{x}}}^{\top}\Sigma^{-1}\partialderivative{\operatorname{\boldsymbol{z}}}{\operatorname{\boldsymbol{x}}}.$
Add Location and Scale parameters for Complex Distributions
In Experiment 5.3 we consider the mixture of two complex distributions. We introduce a location and scale parameter for the components of the mixture. The Funnel, Squiggle and Rosenbrock distributions are smooth bijective transformations from a $Z\sim\operatorname{\mathcal{N}}(\mu,\Sigma)$ to $Y=f(X)$. Let us add a location and scale parameters by an additional transformation $g(y)=x$, where $g(y)=\Sigma_{y}\operatorname{\boldsymbol{I}}+\mu_{y}$
@@
Change of variables for $g\circ f$ gives
@@
Plug in $g^{-1}(x)=\Sigma_{y}^{-1/2}(x-\mu)$, $\det\partialderivative{y}{z}=\det\Sigma^{-1/2}$, and $p_{Z}(z)=\operatorname{\mathcal{N}}(z|\mu,\Sigma)$ to obtain the expression for the density of the Funnel and Squiggle distributions with a location and scale parameters $(\mu_{y},\Sigma_{y})$:
$p_{X}(x)=\operatorname{\mathcal{N}}\left(f^{-1}(\Sigma^{1/2}(x-\mu))\bigg{|}\mu,\Sigma\right)\left|\det\partialderivative{x}{y}\right|\left|\det\Sigma_{y}^{-1/2}\right|.$

### v2

@@ B.2 Numerical integrators
model metr $\mathcal{W}$ min ESS avg ESS avg step-out avg shrinkage t(s)
aus euclidean 0.74 $\pm$ 0.12 18 $\pm$ 5 228 $\pm$ 12 0.14 $\pm$ 0.0 3.33 $\pm$ 0.01 8.2
fisher 0.58 $\pm$ 0.01 177 $\pm$ 9 270 $\pm$ 13 0.95 $\pm$ 0.0 1.0 $\pm$ 0.01 3298.6
ger euclidean 0.49 $\pm$ 0.01 35 $\pm$ 10 119 $\pm$ 9 0.08 $\pm$ 0.0 4.19 $\pm$ 0.02 15.8
monge 0.75 $\pm$ 0.02 80 $\pm$ 5 4673 $\pm$ 113 3.97 $\pm$ 0.03 0.28 $\pm$ 0.01 4665.4
fisher 0.49 $\pm$ 0.0 85 $\pm$ 24 169 $\pm$ 5 0.95 $\pm$ 0.0 0.99 $\pm$ 0.01 19684.4
hrt euclidean 0.63 $\pm$ 0.01 137 $\pm$ 26 254 $\pm$ 16 0.19 $\pm$ 0.01 2.72 $\pm$ 0.02 7.3
monge 0.74 $\pm$ 0.03 394 $\pm$ 55 2979 $\pm$ 266 2.99 $\pm$ 0.04 0.38 $\pm$ 0.02 637.4
fisher 0.64 $\pm$ 0.01 232 $\pm$ 14 311 $\pm$ 8 0.92 $\pm$ 0.01 1.01 $\pm$ 0.01 1694.8
pim euclidean 0.21 $\pm$ 0.0 266 $\pm$ 41 445 $\pm$ 45 0.11 $\pm$ 0.0 3.59 $\pm$ 0.03 6.8
monge 0.29 $\pm$ 0.01 515 $\pm$ 72 4656 $\pm$ 269 2.38 $\pm$ 0.03 0.53 $\pm$ 0.02 1711.7
fisher 0.21 $\pm$ 0.0 427 $\pm$ 47 547 $\pm$ 30 0.93 $\pm$ 0.0 0.98 $\pm$ 0.01 425.4
rip euclidean 0.09 $\pm$ 0.01 829 $\pm$ 112 1499 $\pm$ 56 0.24 $\pm$ 0.01 2.46 $\pm$ 0.02 4.2
monge 0.15 $\pm$ 0.02 1042 $\pm$ 211 3123 $\pm$ 685 1.38 $\pm$ 0.02 0.97 $\pm$ 0.01 593.9
fisher 0.09 $\pm$ 0.0 1623 $\pm$ 113 1753 $\pm$ 92 0.91 $\pm$ 0.0 0.96 $\pm$ 0.01 204.7
Table 1: Bayesian Logistic Regression. Entries are reported as $\mathrm{mean}\pm\mathrm{std}$.
@@ Appendix C Mathematical Derivations
C.1 The Generative and Inverse Generative metrics
The Generative and Inverse Generative metrics read
$G(x)=f(x)I=\exp(\log f(x))I,$
where the scalar factor is $f(x)=\left(\frac{p_{0}+\lambda}{p(x)+\lambda}\right)^{2}$ for the Generative metric and $f(x)=\left(\frac{p(x)+\lambda}{p_{0}+\lambda}\right)^{2}$ for the Inverse Generative metric.
@@ C.1 The Generative and Inverse Generative metrics
The quantities are given by:
$G^{\tfrac{1}{2}}(x)$ $=\exp{\tfrac{1}{2}\log f(x)}I,$
$G^{-\tfrac{1}{2}}(x)$ $=\exp{-\tfrac{1}{2}\log f(x)}I,$
$\log|\det G(\operatorname{\boldsymbol{x}})|$ $=D\log f(x).$
@@
Given the Riemannian metric $G(x)=f(x)I$, the tensor entries are:
@@
The Christoffel symbols for this metric are given by:
@@
Denote by $e_{k}$ the standard basis vectors, the Christoffel symbols in matrix notation are:
$\Gamma^{k}=\frac{1}{2}\left(\nabla\log f(x)\,e_{k}^{\top}+e_{k}\,\nabla\log f(x)^{\top}-\partial_{k}\log f(x)I\right).$
We compute $\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}}=\sum_{i,j=1}^{D}\Gamma^{k}_{ij}(\operatorname{\boldsymbol{x}})\operatorname{\boldsymbol{v}}_{i}\operatorname{\boldsymbol{v}}_{j}$ which appears in the geodesic equations,
@@
The geodesic equations read:
@@
C.2 The Monge and Inverse Monge metrics
The Monge metric and the Inverse Monge metric are:
$G(x)=I+\alpha^{2}\nabla\ell\nabla\ell^{\top},\quad G^{-1}(x)=I-\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell}^{2}}\nabla\ell\nabla\ell^{\top}.$
@@ C.2 The Monge and Inverse Monge metrics
Define the quantity $L_{\alpha}:=1+\alpha^{2}\norm{\nabla\ell}^{2}$, we list the quantities derived from the matrix and present later their derivation,
$G^{1/2}(x)$ $=I+\frac{\alpha^{2}}{1+\sqrt{L_{\alpha}}}\nabla\ell(\operatorname{\boldsymbol{x}})\nabla\ell(\operatorname{\boldsymbol{x}})^{\top},$
$G^{-1/2}(x)$ $=I-\frac{\alpha^{2}}{L_{\alpha}+\sqrt{L_{\alpha}}}\nabla\ell(x)\nabla\ell(x)^{\top},$
$\log|\det G(\operatorname{\boldsymbol{x}})|$ $=\log(1+\alpha^{2}\norm{\nabla\ell}^{2}).$
We now derive the quantities. Hartmann et al. [2022] gave
@@
$G^{-\tfrac{1}{2}}(x)=I+\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1}{\sqrt{L_{\alpha}}}-1\right)\nabla\ell\nabla\ell^{\top}$ (3)
Note that if $\norm{\nabla\ell}^{2}\to 0$ then Equation 3 is undefined. We find a more numerical stable form of $G^{-\tfrac{1}{2}}(x)$ Multiply the scalar $\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1}{\sqrt{L_{\alpha}}}-1\right)$ by its conjugate
$\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1}{\sqrt{L_{\alpha}}}-1\right)=\frac{1}{\norm{\nabla\ell}^{2}}\left(\frac{1-\sqrt{L_{\alpha}}}{\sqrt{L_{\alpha}}}\right)\left(\frac{1+\sqrt{L_{\alpha}}}{1+\sqrt{L_{\alpha}}}\right)=\frac{-\alpha^{2}}{L_{\alpha}+\sqrt{L_{\alpha}}}.$
@@
Plugging the scalar $\frac{-\alpha^{2}}{L_{\alpha}+\sqrt{L_{\alpha}}}$ into $G^{-\tfrac{1}{2}}(\operatorname{\boldsymbol{x}})$, then it is numerically stable for $\norm{\nabla\ell}^{2}\to 0$.
The computation of $G^{\tfrac{1}{2}}(x)$
For convenience take $y=\nabla\ell(x)$. The the metric is $G(y)=I+yy^{\top}$. Let us assume the square root is of the form $G^{\tfrac{1}{2}}(y)=I+\lambda yy^{\top}$. Let us formulate a quadratic equation for $\lambda$:
$G^{\tfrac{1}{2}}(y)G^{\tfrac{1}{2}}(y)$ $=I+yy^{\top}$
$I+2\lambda yy^{\top}+\lambda^{2}\norm{y}^{2}yy^{\top}$ $=I+yy^{\top}$
00 0 $=\left(1-2\lambda-\lambda^{2}\norm{y}^{2}\right)yy^{\top}.$
The solutions of the quadratic equation are
$\lambda=\frac{-1\pm\sqrt{1+\|y\|^{2}}}{\|y\|^{2}}.$
Let us simplify $\frac{-1+\sqrt{1+\|y\|^{2}}}{\|y\|^{2}}$, multiply by its conjugate
$\frac{\sqrt{1+\|y\|^{2}}-1}{\|y\|^{2}}\left(\frac{\sqrt{1+\|y\|^{2}}+1}{\sqrt{1+\|y\|^{2}}+1}\right)=\frac{\norm{y}^{2}}{\norm{y}^{2}\sqrt{1+\|y\|^{2}}+1}.$
Substitute $y=\alpha\nabla\ell(\operatorname{\boldsymbol{x}})$ and we obtain the result
$G^{1/2}(x)=I+\frac{\alpha^{2}}{\sqrt{1+\alpha^{2}\|\nabla\ell(\operatorname{\boldsymbol{x}})\|^{2}}+1}\nabla\ell\nabla\ell^{\top}.$
Christoffel symbols of the Monge metric
The Christoffel associated to the Monge metric (derivation in Section C.3 and Hartmann et al. [2022]) are
$\Gamma^{k}(x)$ $=\frac{\alpha^{2}}{1+\alpha^{2}\norm{\nabla\ell}^{2}}\nabla^{2}\ell\partial_{k}\ell,$
and the geodesic equations read
@@
Christoffel symbols of the Inverse Monge metric
The Christoffel symbols associated to the inverse Monge metric (derivation in Section C.4) are
$\Gamma^{k}=\frac{\alpha^{2}}{2}\Big{[}L_{\alpha}\left(\nabla f\nabla\ell^{\top}+\nabla\ell\nabla f^{\top}+2f\nabla^{2}\ell\right)\partial_{k}\ell+\nabla\ell\nabla\ell^{\top}\partial_{k}f+\alpha^{2}\left\langle\nabla\ell,\nabla f\right\rangle\nabla\ell\nabla\ell^{\top}\partial_{k}\ell\Big{]},$
and the geodesic equations read
@@ C.3 Monge metric: Christoffel symbols derivation
For completeness let us do an alternative derivation of the Christoffel symbols from the one found in Hartmann et al. [2022]. Take the auxiliary function $f(x)=-\frac{1}{L_{\alpha}}$, where $L_{\alpha}=1+\alpha^{2}\norm{\nabla\ell}^{2}$. The metric and inverse components are:
@@
Let us compute $\norm{\operatorname{\boldsymbol{v}}}^{2}_{\Gamma^{k}}$, which appears in the geodesic equations,
@@ C.4 Inverse Monge metric: Christoffel symbols derivation
Again the auxilary function is $f(x)=-\frac{1}{L_{\alpha}}$, where $L_{\alpha}=1+\alpha^{2}\norm{\nabla\ell}^{2}$, the metric and inverse components are
@@
The derivatives of the metric are (we mark with the same color repeating terms)
@@
Let us compute the Christoffel symbols of the first kind (blue and red terms will cancel out, pink terms add to each other)
@@
The Christoffel symbols of the second kind read
@@
Written in matrix form
@@
Let us compute $\norm{v}^{2}_{\Gamma^{k}}$ which appears in the geodesic equations
@@
Then the geodesic equations read,
@@
C.5 Target Distributions
The Funnel, Squiggle and Rosenbrock distributions are smooth bijective transformations from a $Z\sim\operatorname{\mathcal{N}}(\mu,\Sigma)$ to $X=f(X)$. We use the shorthand notation $x=x(z)$ and $z=z(x)$.
@@ C.5 Target Distributions
$p(x)=\mathcal{N}(x_{D}|0,\sigma^{2})\mathcal{N}(x_{1:{D-1}}|\mu,e^{x_{D}}I_{D-1})$. In this case $Z\sim\operatorname{\mathcal{N}}(0,I)$. The choice of parameters is $\sigma=3$ and $\mu=0$,
@@
For simplicity here we show the two dimensional case, the full distribution can be consulted in Pagani et al. [2022]. The two dimensional density is: $p(x)=\operatorname{\mathcal{N}}(x_{1}|a,\frac{1}{2})\operatorname{\mathcal{N}}(x_{2}|x_{1}^{2},\frac{1}{2b})$. In this case $Z\sim\operatorname{\mathcal{N}}(0,I)$. The choice of parameters is $a=1$, $b=100$ and block size of $3$ and $\lfloor\tfrac{D-1}{3}\rfloor$ total blocks,
@@
The density is $p(x)=\operatorname{\mathcal{N}}(x(z)|\mu,\Sigma)|\det\partialderivative{x}{z}|$, where $Z\sim\operatorname{\mathcal{N}}(\mu,\Sigma)$. The choice of parameters is $a=1.5$, $\mu=0$, $\Sigma=\operatorname{\mathrm{diag}}(5,\tfrac{1}{2},..,\tfrac{1}{2})$
@@
For these three toy problems the Fisher Information follows from the transformation rule of Riemannian metrics
$G(x)=\partialderivative{z}{x}^{\top}\Sigma^{-1}\partialderivative{z}{x}.$
Location and Scale parameters for Complex Distributions
In Experiment 5.3 we consider the mixture of two complex distributions. We introduce a location and scale parameter for each components of the mixture. The Funnel, Squiggle and Rosenbrock distributions are smooth bijective transformations from a $Z\sim\operatorname{\mathcal{N}}(\mu,\Sigma)$ to $Y=f(X)$. Let us add a location and scale parameters by an additional transformation $g(Y)=X$, where $g(y)=\Sigma_{y}y+\mu_{y}$
@@
The change of variable formula for the composition $g\circ f$ gives
@@
Plug in $g^{-1}(x)=\Sigma_{y}^{-1/2}(x-\mu_{y})$, $\det\partialderivative{y}{z}=\det\Sigma^{-1/2}$, and $p_{Z}(z)=\operatorname{\mathcal{N}}(z|\mu,\Sigma)$, we obtain the expression of the density
$p_{X}(x)=\operatorname{\mathcal{N}}\left(f^{-1}(\Sigma^{1/2}(x-\mu_{y}))\bigg{|}\mu,\Sigma\right)\left|\det\partialderivative{x}{y}\right|\left|\det\Sigma_{y}^{-1/2}\right|.$
Where $(\mu_{y},\Sigma_{y})$ are the location and scale parameters of the component of the mixture distribution.
The Allen-Cahn Field System
We consider the stochastic Allen–Cahn model [Berglund et al., 2017] used as a benchmark in Cabezas et al. [2024b]. The log-density is:
$\log p(x)=-\beta\left(\frac{a}{2\Delta s}\sum_{i=1}^{D+1}(x_{i}-x_{i-1})^{2}+\frac{b\Delta s}{4}\sum_{i=1}^{D}(1-x_{i}^{2})^{2}\right).$ (4)
We adopt the parameter choices $\Delta s=\tfrac{1}{D}$ and boundary conditions $x_{0}=x_{D+1}=0$, and the constants $a=0.1$ and $b=\tfrac{1}{a}$ ensure that the double-well potential induces bimodality in each component $x_{i}$, and we fix $D=16$.
Analysis of multimodality
To understand the maxima of this density, we analyze the two terms in the log-density function (Eq. 4):
  1. 1.
    The first term, $\sum_{i=1}^{D+1}(x_{i}-x_{i-1})^{2}$, penalizes differences between adjacent components, encouraging all components to have similar values.
  2. 2.
    The second term, $\sum_{i=1}^{D}(1-x_{i}^{2})^{2}$, is minimized when $x_{i}=\pm 1$.
The global maxima occur at $(1,\ldots,1)$ and $(-1,\ldots,-1)$ because these configurations minimize both terms simultaneously: all components have the same value (satisfying the first term) and each component equals ±$\pm 1$ (satisfying the second term).
Local maxima occur at all other combinations of ±$\pm 1$ values (i.e., at points (±$(\pm 1,\ldots,\pm 1)$ with mixed signs) because these configurations still satisfy the second term perfectly, but incur penalties from the first term due to sign changes between adjacent components.
This creates $2^{D}$ local maxima, making the problem highly multimodal, with the two homogeneous configurations being global maxima.
Kernel Stein Discrepancy
Let $\pi$ and $\nu$ be two probability measures. We estimate the Kernel Stein Discrepancy with the biased but non-negative V-estimator. Given a sample $x_{i}\sim\nu$ for $i=1,\ldots,n$,
$\widehat{\mathrm{KSD}}^{2}_{k,V}(\pi,\nu)=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k_{\pi}(x_{i},x_{j}^{\prime}).$
Denote by $p(x)$ be the density of the measure $\pi$, then,
$k_{\pi}(x,x^{\prime})=\nabla_{x}\cdot\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x}k(x,x^{\prime})\cdot\nabla_{x^{\prime}}\log p(x^{\prime})+\nabla_{x^{\prime}}k(x,x^{\prime})\cdot\nabla_{x}\log p(x)+k(x,x^{\prime})\nabla_{x}\log p(x)\cdot\nabla_{x^{\prime}}\log p(x),$
where we choose the inverse multi quadratic kernel $k(x,x^{\prime})=(1+(x-x^{\prime})^{\top}(x-x^{\prime}))^{\beta}$ for $\beta=-\tfrac{1}{2}$, following the choices made by Cabezas et al. [2024b].